OpenStudy (anonymous):

A triangular paddock has a road AB forming its hypoteneuse. AB is 3km long. The fences AC and CB are at right angles. If BC is 400m longer than AC, how do you find the area of the paddock? (in hectares)

OpenStudy (anonymous):

You need to solve teh quadratic \[3000^2=(AC+400)^2+BC^2\] (from Pythagoras' theorem)

OpenStudy (anonymous):

Opps. That last term should be \[AC^2\] also

OpenStudy (anonymous):

9 = x^2 +(x+400)^2 solve for x to get x ~ 1.91187 km then, just multiply 1.91187 by (1.91187 + 0.40000) to get area of paddock ~ 4.41999... km2 then just convert km2 to hectares 1 km2 = 100 ha so 4.419999 km2 = 441.999... hectares ~ 442 hectares

OpenStudy (anonymous):

did I mention how cute you are?

OpenStudy (anonymous):

yes.. and oh my.. i'm flattered.. haha

OpenStudy (anonymous):

<3

OpenStudy (saifoo.khan):

lol.